# Prospects for detecting asteroid-mass primordial black holes in extreme-mass-ratio inspirals with continuous gravitational waves

### v1

@@ V Projected constraints
ET will provide unprecedented low-frequency sensitivity to GWs, enabling us to see much longer signals arising from the inspirals of two compact objects than currently possible with LIGO, Virgo and KAGRA. Furthermore, the sensitivity across all frequencies will increase by approximately an order of magnitude, enabling numerous detections of binary black hole and binary neutron star mergers Maggiore et al. (2020); Branchesi et al. (2023).
It is thus worth asking to what extent ET will be able to detect asteroid-mass PBHs that could form in a binary system. Following the methodology outlined in Section IV, and considering a range of possible EMRI systems, we compute the expected luminosity distance reach and constraints on $\tilde{f}$, as shown in Fig. 2(a). We have selected the maximum distance reach possible for each system to plot here, that is, selecting the particular frequency at which the distance reach is maximum. Additionally, we have assumed $\dot{f}_{\rm max}<10^{-9}$ Hz/s, which was used in Abbott et al. (2022a), and a frequency evolution given by Eq. 4. Though we make these plots in terms of $m_{1}$ and $m_{2}$, the distance reach depends primarily on $\mathcal{M}$, which is constant along the diagonal lines in these plots. When computing $\tilde{f}$, however, we use the contributions for EMRI systems inspiraling at all frequencies, not just the maximum one, as implied by Eq. 14. We evaluate the impact of the range of frequencies chosen on the $\tilde{f}$ constraint in Section VI.
@@ VI.1 Varying $\dot{f}_{\rm max}$
A major limitation of CW searches to probe asteroid-mass PBHs is the linear signal model (Eq. 11) and the $\dot{f}_{\rm max}$ considered. These two criteria are linked: systems with chirp masses above a critical value will require us to use second- or third-order terms in Eq. 4 to model them correctly. Furthermore, even among systems that do follow Eq. 4, $\dot{f}_{\rm max}$ prevents higher frequencies from contributing to the sum in Eq. 14, which, as we will argue, degrades the constraint on $\tilde{f}$.
We thus ask the question: if CW searches retain the signal model in Eq. 4 but could increase $\dot{f}_{\rm max}$, how would the ability to constrain $\tilde{f}$ change in ET? We provide an answer to this question in Fig. 3. These plots show that for $10^{-9}$ Hz/s $\leq\dot{f}_{\rm max}\leq 10^{-7}$ Hz/s does not alter the sensitivity much, since the linearity condition in Eq. 4 is violated. Thus, if we remove both the linearity condition and $\dot{f}_{\rm max}$, which would, in practice, require asteroid-mass PBHs to be searched for with different methods Miller et al. (2021b); Andrés-Carcasona et al. (2023); Alestas et al. (2024), we find orders of magnitude improvement in the constraint on $\tilde{f}$. We will evaluate the suitability of these different methods to probe asteroid-mass PBHs in Section VI.4.
Figure 3: Left: Varying the maximum spin-up to which CW searches are sensitive results in different constraints on $\tilde{f}$. Current searches consider $\dot{f}_{\rm max}=10^{-9}$ Hz/s. We can see that smaller $\dot{f}$ indicates not only a poorer sensitivity at small $m_{2}$, but also cannot reach higher values of $m_{2}$, since the signal will spin up to values higher than $\dot{f}_{\rm max}$ during $T_{\text{obs}}$. The degradation in sensitivity of smaller $\dot{f}_{\rm max}$ occurs because signals at higher frequencies cannot contribute to the sum in Eq. 14, since the signal would either take on $\dot{f}\geq\dot{f}_{\rm max}$, and/or the GW frequency evolution cannot be described by Eq. 4 anymore. Right: A comparison showing how much the constraints would improve if no $\dot{f}_{\rm max}$ existed, that is, if the signal could be searched for at arbitrarily high $\dot{f}$ with a frequency evolution following Eq. 3. In both plots, we have set $\delta f=1$ Hz, which represents the approximate spacing in upper limits that is obtained through injections in CW searches Abbott et al. (2022a).
@@ VI.2 Impact of eccentricity
Since we are considering binary black holes with such extreme mass ratios, we should evaluate to what extent our results are valid if these systems are eccentric. To begin, we evaluate the impact of eccentricity on the PN0 $\dot{f}$ term (Eq. 2) Maggiore (2008):
@@
where $g(\epsilon)$ is a function arising from considering GW emission in a quasi-elliptical orbit of two masses, and $\epsilon$ is the eccentricity of the system.
@@
How eccentric inspiraling systems are depends on how they formed Kowalska et al. (2011); Samsing et al. (2014); Samsing (2018). Additionally, only limited models exist for eccentricity evolution as a function of time Arredondo et al. (2024). We thus assume the “worst-case” scenario in which the eccentricity is constant throughout the orbit, thus causing the maximum shift in $\dot{f}$. In Fig. 4, we show the impact of eccentricity on the projected constraints for $\tilde{f}$ for $m_{1}=2.5M_{\odot}$ using the ET sensitivity curve. We follow the procedure outlined in Section IV to obtain the distance reached and thus the constraint on $\tilde{f}$.
In Fig. 4, we notice that as the eccentricity increases, the maximum $m_{2}$ probable with CW methods decreases. This is because $\dot{f}_{\rm ecc}$ becomes too large and thus the condition in Eq. 19 is no longer satisfied. Furthermore, we see a general weakening of the constraint on $\tilde{f}$ as we increase eccentricity, since the systems with higher frequencies no longer contribute to the sum because the condition in Eq. 19 is more easily violated at high frequencies even though it is met at low frequencies.
Figure 4: ET. We vary the eccentricity and compute the expected constraint on $\tilde{f}$ for a fixed $m_{1}=2.5M_{\odot}$. Highly eccentric systems are much harder to constrain with standard CW searches.
@@ VI.3 From which frequencies does the sensitivity come?
Ironically, the restrictions on linearity in current CW searches imply that lower-frequency signals will always contribute to the sum, but they do not comprise a large gain in sensitivity with respect to those at high frequencies, thus motivating the need to attempt to detect signals at high frequencies and removing restrictions on $\dot{f}_{\rm max}$ and linearity. We thus study specifically which distance reaches, indexed by the GW frequencies in the sum in Eq. 14, contribute the most to the rate density constraint. We show in Fig. 5(a) constraints on the rate density for particular choices of $m_{1}$ and $m_{2}$, assuming negligible eccentricity, a linear frequency evolution (Eq. 4) and $\dot{f}_{\rm max}=10^{-9}$ Hz/s. The way to intepret this plot is as follows: at a given frequency, the sum over distance reaches is taken from that frequency to the maximum one in the plot. For example, for the red curve, at $f_{\text{min}}\simeq 200$ Hz, we sum the distance reaches from $f_{\text{min}}\simeq 200$ Hz to $f_{\text{max}}=800$ Hz. We notice immediately in the red curve that going to frequencies below 200 Hz does not improve the constraint on the rate density, indicating that the constraint on $\tilde{f}$ does not get better by looking at lower GW frequencies. We can see this behavior in each of the curves on this plot, except that the “cut-off” frequency below which it is no longer is useful to sum the distance reaches in Eq. 14 decreases. This occurs because as $m_{2}$ increases, for a fixed $m_{1}$, $\mathcal{M}$ increases and thus $\dot{f}$ increases, indicating that, for higher frequencies, the GW frequency evolution will deviate from Eq. 4. We can see that for the highest value of $m_{2}$, that summing the distance reaches at all available frequencies is immensely helpful. Additionally, Fig. 5 shows the frequency range $[f_{\text{min}},f_{\text{max}}]$ necessary to analyze to obtain 99% of the sensitivity on the rate density value, as a function of $m_{2}$. This plot is consistent with Fig. 5(a), showing that for smaller values of $m_{2}$, higher frequencies provide the best sensitivity.
Figure 5: ET sensitivity curve. $m_{1}=2.5M_{\odot}$. Left: Accumulated rate density constraint as a function at which frequency we begin to sum the distance reaches in Eq. 14 up to a maximum given by the right-most frequency on each curve. Right: the frequency band to analyze to obtain 99% of the optimal rate density constraint as a function of $m_{2}$, where the “optimal” constraint would arise if we sum distance reaches over the full frequency range.
@@ VI.4 Changes to current CW searches
As an example of what could be achieved in the previous observing run of advanced LIGO, Virgo and KAGRA (O3), we show in Fig. 6 how the constraints on $\tilde{f}$ in Abbott et al. (2022a) would change if $\dot{f}_{\rm max}=10^{-9}$ Hz/s were increased and if the linearity condition in Eq. 4 were relaxed. In Abbott et al. (2022a), $\delta f=\dot{f}_{\rm max}T_{\text{obs}}$ [^2] This choice implies that, at each frequency at which an upper limit is set, that the signal frequency cannot vary by more than $\delta f$ over the course of $T_{\text{obs}}$., so the spacing in the O3 interpolated upper limits changes when computing $\tilde{f}$ changes as a function of $\dot{f}_{\rm max}$. Allowing $\dot{f}_{\rm max}\rightarrow 100\dot{f}_{\rm max}$ results in weaker constraints on $\tilde{f}$ than both the magenta and blue curves at certain masses. This occurs because, as shown in Appendix B, there is a non-trivial dependence of $\tilde{f}$ on the spacing in frequency $\delta f$ in the case of a flat power spectral density, in which larger $\delta f$ implies worse limits. However, the actual dependence of $\tilde{f}$ on $\delta f$, i.e. when including a frequency-dependent power spectral density and conditions on linearity and $\dot{f}_{\rm max}$, result in a complicated behavior of these constraints, shown in Fig. 6.
Figure 6: How the constraint on $\tilde{f}$ obtained in Abbott et al. (2022a) would change if various conditions of CW searches were relaxed.
@@
Fig. 6 implies that, even now, constraints of $\tilde{f}\lesssim 1$ could be obtained at masses below $\sim 10^{-10}M_{\odot}$; however, relaxing the requirement in Eq. 4 and allowing $100\dot{f}_{\rm max}=10^{-7}$ Hz/s has implications for the computational cost of the search. Increasing $\dot{f}_{\rm max}$ would imply an increase in the size of the grid in $\dot{f}$ that is analyzed in CW searches. Currently, the step-size in $\delta\dot{f}=1/(T_{\text{FFT}}T_{\text{obs}})\sim 3\times 10^{-11}$ Hz/s at high frequencies Abbott et al. (2022a), where $T_{\text{FFT}}=1024$ s. To extend the positive $\dot{f}$ coverage to $\dot{f}_{\rm max}=10^{-8}$ Hz/s and $\dot{f}_{\rm max}=10^{-7}$ Hz/s would require 3300 and 33000 points more to search over, respectively, which is a 10x to 100x increase in the the number of points in the current spin-down grid, which are $\sim 320$ negative values and $\sim 32$ positive ones, at each sky position. Such an increase in computational cost is not currently feasible, since all-sky CW searches already take $10^{7}$ CPU core-hours Abbott et al. (2022a); Miller et al. (2024b), and would not lead to a powerful constraint on $\tilde{f}$ at current detector sensitivity unless the linearity condition was relaxed, which would, again, imply increased computational cost to consider $\ddot{f}$ or $\dddot{f}$ terms in Eq. 4. However, searching at a specific point in the sky, e.g. the Galactic Center, would be feasible, since the current “directed” searches towards particular sky positions have orders of magnitude less computational cost than the all-sky ones, since the Doppler shift induced by the relative motion of the earth with respect to the source can be corrected for before doing the analysis Abbott et al. (2022c).
It is thus worth asking if other methods could be used to achieve the constraints given by the black curve now, that is to search for signals that do not strictly follow Eq. 4. Such methods, which are termed tCW methods, search directly for Eq. 3, and thus are sensitive to arbitrarily high values of $\dot{f}$.
As an example, we evaluate the computational cost of the Generalized frequency-Hough, a particular tCW method that finds curves in the time-frequency plane Miller et al. (2018, 2021b, 2024b), at a fixed sky position, for $T_{\text{FFT}}=8192$ s. In this method, a grid in $k$, instead of $\dot{f}$, is constructed, and the loop over each point in this grid constitutes the greatest computational cost of the search.
The spacing in the grid $dk$ in our case is Miller et al. (2018):
$dk\approx\frac{11}{3}\left(\frac{k}{T_{\text{FFT}}f_{\text{gw}}}\right)$ (20)
and varies between $[k_{\rm min},k_{\rm max}]$, where $k_{\rm min}=\dot{f}_{\rm min}/f_{\text{max}}$ and $k_{\rm max}=\dot{f}_{\rm max}/f_{\text{min}}$. The number of points in the $k$ grid is therefore a function of both the specified $f$ and $\dot{f}$ range. Furthermore, instead of frequency, there is another grid that has spacing $dx$:
$dx=\frac{8}{3}\frac{1}{T_{\text{FFT}}f_{\text{gw}}^{11/3}}$ (21)
which ranges from $x_{\rm min}=1/f_{\text{max}}^{8/3}$ and $x_{\rm max}=1/f_{\text{min}}^{8/3}$.
In Fig. 7, we show how the number of points in each grid changes as a function of each 5-Hz frequency band analyzed. In total, knowing that each $x/k$ point takes about $50\mu$s Miller et al. (2021b) to evaluate with the Generalized frequency-Hough (though this time is likely to decrease by an order of magnitude or even more [^3] Private correspondence with Lorenzo Pierini), we calculate that a search using the Generalized frequency-Hough at one sky position would take $\sim 2.8\times 10^{4}$ core-hours on one Quad-Core Intel Core i7. Such a search of one sky position is orders of magnitude cheaper than those presented in Abbott et al. (2022a).
Figure 7: The number of points in the two grids, $x$ and $k$, that relate to $f_{\text{gw}}$ and $\dot{f}$, respectively, that would be required to search over using the Generalized frequency-Hough. We consider frequency bands of 5 Hz, a spin-up range of $[10^{-9},10^{-7}]$ Hz/s, and $T_{\text{FFT}}=8192$ s Abbott et al. (2022a), since we restrict ourselves low frequencies ($<100$ Hz) to facilitate direct comparisons with Fig. 5 and because the computational cost increases with $f_{\text{gw}}$.
We can also evaluate how many sky points, $N_{\text{sky}}$, would be present at low frequencies in order to actually perform an “all-sky search” for such asteroid-mass PBH systems using data from current detectors. Here, we focus $f_{\text{gw}}<200$ because for the values of $m_{2}$ in Fig. 6 at which we would hope to have a constraint on $\tilde{f}$ in current detector data, these low frequencies contribute the most to the sum in Eq. 14, as shown in Fig. 8 and demonstrated earlier in Fig. 5, and because $N_{\text{sky}}$ scales with the square of the frequency and $T_{\text{FFT}}$:
$N_{\text{sky}}\simeq 4\pi T_{\text{FFT}}^{2}f^{2}\frac{v_{\rm orb}^{2}}{c^{2}}$ (22)
where $v_{\rm orb}$ is the velocity of earth around the solar-system barycenter. We plot the number of sky points as a function of frequency in Fig. 9 for $T_{\text{FFT}}=8192$ s and for $T_{\text{FFT}}=1.5$ days. With the expected improvement in the Generalized frequency-Hough of about a factor of 10, searching up to $\sim 200$ Hz would require $\sim 2.8\times 10^{8}$ core-hours, which is a little more than current all-sky searches up to 2048 Hz cost. If we decrease the maximum frequency range to, say, 100 Hz, we would obtain roughly a factor of $\sim 3-4$ decrease in computational cost, and still be sensitive to asteroid-mass PBHs.
Figure 8: Distance reach as a function of frequency for the “$\dot{f}=100\dot{f}_{\rm max}$, no linear approximation” case shown in Fig. 6 (solid black curve in that figure), restricted to showing a few values of $m_{2}$ (different colors) for which $\tilde{f}\lesssim 1$ in O3.
Figure 9: Number of sky points as a function of GW frequency for two choices of $T_{\text{FFT}}$. The green curve corresponds to the $T_{\text{FFT}}$ used in current all-sky searches for CWs, while the blue curve is drawn with a much longer $T_{\text{FFT}}$ that we expect to use in the ET era.

### v2

@@ V Projected constraints
ET will provide unprecedented low-frequency sensitivity to GWs, enabling us to see much longer signals arising from the inspirals of two compact objects than currently possible with LIGO, Virgo and KAGRA. Furthermore, the sensitivity across all frequencies will increase by approximately an order of magnitude, enabling numerous detections of binary black hole and binary neutron star mergers Maggiore:2019uih; Branchesi:2023mws.
It is thus worth asking to what extent ET will be able to detect asteroid-mass PBHs that could form in a binary system. Following the methodology outlined in Section IV, and considering a range of possible EMRI systems, we compute the expected luminosity distance reach and constraints on $\tilde{f}$, as shown in Fig. 2(a). We have selected the maximum distance reach possible for each system to plot here, that is, selecting the particular frequency at which the distance reach is maximum. Additionally, we have assumed $\dot{f}_{\rm max}<10^{-9}$ Hz/s, which was used in KAGRA:2022dwb, and a frequency evolution given by Eq. 4. Though we make these plots in terms of $m_{1}$ and $m_{2}$, the distance reach depends primarily on $\mathcal{M}$, which is constant along the diagonal lines in these plots. When computing $\tilde{f}$, however, we use the contributions for EMRI systems inspiraling at all frequencies, not just the maximum one, as implied by Eq. 15. We evaluate the impact of the range of frequencies chosen on the $\tilde{f}$ constraint in Section VI.
@@ VI.1 Varying $\dot{f}_{\rm max}$
A major limitation of CW searches to probe asteroid-mass PBHs is the linear signal model (Eq. 12) and the $\dot{f}_{\rm max}$ considered. These two criteria are linked: systems with chirp masses above a critical value will require us to use second- or third-order terms in Eq. 4 to model them correctly. Furthermore, even among systems that do follow Eq. 4, $\dot{f}_{\rm max}$ prevents higher frequencies from contributing to the sum in Eq. 15, which, as we will argue, degrades the constraint on $\tilde{f}$. Essentially, through (2), the maximum $\dot{f}$, for a given chirp masses, fixes the maximum frequency that we can use to constrain $\tilde{f}$.
We thus ask the question: if CW searches retain the signal model in Eq. 4 but could increase $\dot{f}_{\rm max}$, how would the ability to constrain $\tilde{f}$ change in ET? We provide an answer to this question in Fig. 3. These plots show that for $10^{-9}$ Hz/s $\leq\dot{f}_{\rm max}\leq 10^{-7}$ Hz/s, the projected constraints on $\tilde{f}$ are almost identical, since the linearity condition in Eq. 4 is violated. Thus, if we remove both the linearity condition and $\dot{f}_{\rm max}$, which would, in practice, require asteroid-mass PBHs to be searched for with different methods Miller:2020kmv; Andrés-Carcasona:2023df; Alestas:2024ubs, we find orders of magnitude improvement in the constraint on $\tilde{f}$. We will evaluate the suitability of these different methods to probe asteroid-mass PBHs in Section VI.4.
Figure 3: Left: Varying the maximum spin-up to which CW searches are sensitive results in different constraints on $\tilde{f}$. Current searches consider $\dot{f}_{\rm max}=10^{-9}$ Hz/s. We can see that smaller $\dot{f}$ indicates not only a poorer sensitivity at small $m_{2}$, but also an inability to reach higher values of $m_{2}$, since the signal spin up will increase to be higher than $\dot{f}_{\rm max}$ during $T_{\text{obs}}$. The degradation in sensitivity of smaller $\dot{f}_{\rm max}$ occurs because signals at higher frequencies cannot contribute to the sum in Eq. 15, since the signal would either take on $\dot{f}\geq\dot{f}_{\rm max}$, and/or the GW frequency evolution cannot be described by Eq. 4 anymore. Right: A comparison showing how much the constraints would improve if no $\dot{f}_{\rm max}$ existed, that is, if the signal could be searched for at arbitrarily high $\dot{f}$ with a frequency evolution following Eq. 3. In both plots, we have set $\delta f=1$ Hz, which represents the approximate spacing in upper limits that is obtained through injections in CW searches KAGRA:2022dwb.
@@ VI.2 Impact of eccentricity
Since we are considering binary black holes with such extreme mass ratios, we should evaluate to what extent our results are valid if these systems are eccentric. To begin, we evaluate the impact of eccentricity on the PN0 $\dot{f}$ term (Eq. 2) maggiore2008gravitational:
@@
where $g(\epsilon)$ is a function arising from considering GW emission in a quasi-elliptical orbit of two masses, and $\epsilon$ is the eccentricity of the system. The above equations describe how GWs are emitted more quickly with respect to those that arise from circular orbits, and come from an average over each orbit Peters:1963ux; Peters:1964zz. They cannot account for any instantaneous variations in the time-frequency evolution Amaro-Seoane:2007osp, such as harmonics Teukolsky:1973ha; Drasco:2005kz; Fujita:2009bp; Katz:2021yft, bursts Glampedakis:2002cb; Rubbo:2006vh; Hopman:2006fc; Toonen:2009qi; Berry:2013poa; Oliver:2023xan, transient resonances Pound:2007th; Barack:2009ux; Flanagan:2010cd; Gair:2010iv; Poisson:2011nh; Flanagan:2012kg; Berry:2016bit, etc. In other words, these instantaneous frequency variations must be confined to one frequency bin in each $T_{\text{FFT}}$, and there can be no large-scale shift in the time-frequency evolution that deviates from Eq. 4. We thus only provide projected constraints in the regime in which such instantaneous, non-monotonic variations would be contained within a frequency bin and/or do not occur.
@@
How eccentric inspiraling systems are depends on how they formed Kowalska:2010qg; Samsing:2013kua; Samsing:2017xmd. Additionally, only limited models exist for eccentricity evolution as a function of time Arredondo:2024nsl. We thus assume the “worst-case” scenario in which the eccentricity is constant throughout the orbit, thus causing the maximum shift in $\dot{f}$. In Fig. 4, we show the impact of eccentricity on the projected constraints for $\tilde{f}$ for $m_{1}=2.5M_{\odot}$ using the ET sensitivity curve and assuming $\dot{f}_{\rm max}=10^{-9}$ Hz/s. We follow the procedure outlined in Section IV to obtain the distance reached and thus the constraint on $\tilde{f}$.
In Fig. 4, we notice that as the eccentricity increases, the maximum $m_{2}$ probeable with CW methods decreases. This is because $\dot{f}_{\rm ecc}$ becomes too large and thus the condition in Eq. 20 is no longer satisfied. Furthermore, we see a general weakening of the constraint on $\tilde{f}$ as we increase eccentricity, since the systems with higher frequencies no longer contribute to the sum because the condition in Eq. 20 is more easily violated at high frequencies even though it is met at low frequencies.
Figure 4: We vary the eccentricity and compute the expected constraint on $\tilde{f}$ for a fixed $m_{1}=2.5M_{\odot}$ for a search that assumes $\dot{f}_{\rm max}=10^{-9}$ Hz/s and using the ET power spectral density. Highly eccentric systems are much harder to constrain with standard CW searches than lower ones, $\dot{f}_{\rm ecc}$ exceeds $\dot{f}_{\rm max}$ much more easily than for low-eccentricity systems.
@@ VI.3 From which frequencies does the sensitivity come?
Ironically, the restrictions on linearity in current CW searches imply that lower-frequency signals will always contribute to the sum, but they do not comprise a large gain in sensitivity with respect to those at high frequencies, thus motivating the need to attempt to detect signals at high frequencies and removing restrictions on $\dot{f}_{\rm max}$ and linearity. We thus study specifically which distance reaches, indexed by the GW frequencies in the sum in Eq. 15, contribute the most to the rate density constraint. We show in Fig. 5 constraints on the rate density for particular choices of $m_{1}$ and $m_{2}$, assuming negligible eccentricity, a linear frequency evolution (Eq. 4) and $\dot{f}_{\rm max}=10^{-9}$ Hz/s. The way to interpret this plot is as follows: at a given frequency, the sum over distance reaches is taken from that frequency to the maximum one in the plot. For example, for the red curve, at $f_{\text{min}}\simeq 200$ Hz, we sum the distance reaches from $f_{\text{min}}\simeq 200$ Hz to $f_{\text{max}}=800$ Hz. We notice immediately in the red curve that going to frequencies below 200 Hz does not improve the constraint on the rate density, indicating that the constraint on $\tilde{f}$ does not get better by looking at lower GW frequencies. We can see this behavior in each of the curves on this plot, except that the “cut-off” frequency below which it is no longer is useful to sum the distance reaches in Eq. 15 decreases. This occurs because as $m_{2}$ increases, for a fixed $m_{1}$, $\mathcal{M}$ increases and thus $\dot{f}$ increases, indicating that, for higher frequencies, the GW frequency evolution will deviate from Eq. 4. We can see that for the highest value of $m_{2}$, that summing the distance reaches at all available frequencies is immensely helpful.
Figure 5: Accumulated rate density constraint as a function of which frequency we begin to sum the distance reaches in Eq. 15 up to a maximum given by the right-most frequency on each curve. This has been computed using the ET power spectral density and assuming $m_{1}=2.5M_{\odot}$. The frequency on the x-axis represents the minimum frequency considered in the range $f_{\text{min}},f_{\text{max}}]$. For all values of $m_{2}$ plotted here, it is clear that the rate density saturates at a particular minimum frequency, which means that it is no longer beneficial to consider frequencies below $f_{\text{min}}$ in the sum to compute constraints on the rate density.
@@ VI.4 Changes to current CW searches
As an example of what could be achieved in the previous observing run of advanced LIGO, Virgo and KAGRA (O3), we show in Fig. 6 how the constraints on $\tilde{f}$ in KAGRA:2022dwb would change if $\dot{f}_{\rm max}=10^{-9}$ Hz/s were increased and if the linearity condition in Eq. 4 were relaxed. In KAGRA:2022dwb, $\delta f=\dot{f}_{\rm max}T_{\text{obs}}$ is used [^3] This choice implies that, at each frequency at which an upper limit is set, that the signal frequency cannot vary by more than $\delta f$ over the course of $T_{\text{obs}}$., so the spacing in the O3 interpolated upper limits changes when computing $\tilde{f}$ changes as a function of $\dot{f}_{\rm max}$. Allowing $\dot{f}_{\rm max}\rightarrow 100\dot{f}_{\rm max}$ results in weaker constraints on $\tilde{f}$ than both the magenta and blue curves at certain masses. This occurs because, as shown in Appendix A, there is a non-trivial dependence of $\tilde{f}$ on the spacing in frequency $\delta f$ in the case of a flat power spectral density, in which larger $\delta f$ implies worse limits. However, the actual dependence of $\tilde{f}$ on $\delta f$, i.e. when including a frequency-dependent power spectral density and conditions on linearity and $\dot{f}_{\rm max}$, result in a complicated behavior of these constraints, shown in Fig. 6.
Figure 6: How the constraint on $\tilde{f}$ obtained in KAGRA:2022dwb would change if various conditions of CW searches were relaxed. In particular, we compare results from O3 to the following conditions: (1) remove the requirement that the time-frequency evolution be linear; (2) increase $\dot{f}_{\rm max}$ by a factor of 10 and 100, but still require the time-frequency evolution to be linear; and (3) again increase $\dot{f}_{\rm max}$ by a factor of 10 and 100, but relax the requirement the time-frequency evolution to be linear. We can see that, even with data from the current observing run, if the CW methods could handle nonlinear time-frequency evolutions and also search up to a higher $100\dot{f}_{\rm max}$, we would be able to constrain $\tilde{f}<1$ for $m_{2}\in[3\times 10^{-10},10^{-7}]M_{\odot}$.
@@
Fig. 6 implies that, even now, constraints of $\tilde{f}\lesssim 1$ could be obtained at masses below $\sim 10^{-10}M_{\odot}$; however, relaxing the requirement in Eq. 4 and allowing $100\dot{f}_{\rm max}=10^{-7}$ Hz/s has implications for the computational cost of the search. Increasing $\dot{f}_{\rm max}$ would imply an increase in the size of the grid in $\dot{f}$ that is analyzed in CW searches. Currently, the step size is $\delta\dot{f}=1/(T_{\text{FFT}}T_{\text{obs}})\sim 3\times 10^{-11}$ Hz/s at high frequencies KAGRA:2022dwb, where $T_{\text{FFT}}=1024$ s. To extend the positive $\dot{f}$ coverage to $\dot{f}_{\rm max}=10^{-8}$ Hz/s and $\dot{f}_{\rm max}=10^{-7}$ Hz/s would require an additional 3,300 and 33,000 steps to be searched, respectively. This represents a 10x to 100x increase in the number of steps in spin-down, which currently includes approximately 320 negative steps and 32 positive steps at each sky position, which results in 10-100x more points in the parameter space to analyze. Such an increase in computational cost is not currently feasible, since all-sky CW searches already take $10^{7}$ CPU core-hours KAGRA:2022dwb; Miller:2024jpo, and would not lead to a powerful constraint on $\tilde{f}$ at current detector sensitivity unless the linearity condition was relaxed, which would, again, imply increased computational cost to consider $\ddot{f}$ or $\dddot{f}$ terms in Eq. 4. However, searches that target a specific point in the sky, e.g. the Galactic Center, could expand the range and number of spin-down parameters that are analyzed in order in order to probe less linear regimes for these signals.